# Denoising diffusion probabilistic models are optimally adaptive to unknown low dimensionality

### v2

@@ 5 Main analysis (Proof of Theorem 2)
We now present the proof of main convergence theory in Theorem 2. To control the KL divergence between the DDPM-type dynamics and our target distribution, a common approach is to connect the KL error with the $\ell_{2}$ discretization error by means of Girsanov’s theorem (e.g., Chen et al., 2022b ; Chen et al., 2023b ). We shall follow this approach in this proof, while in the meantime taking into account the intrinsic low-dimensional structure of the target distribution. In particular, we shall invoke the covering argument to treat the low-dimensional structure approximately as the union of linear subspaces. In particular, As it turns out, the evolution of the posterior mean function, i.e., $\mu_{t}(\cdot)$, can be well calibrated along the forward process with high probability under this framework, which plays a key role in enabling our non-asymptotic convergence analysis.
Our main proof is composed of three steps, with the first two steps being fairly standard and the third step capturing the main challenges. The analysis is supported by a sequence of auxiliary lemmas whose proofs are deferred to Appendices B and C. Throughout the proof, we adopt the following convenient notation:
@@
  • $P^{T}$: the measure of the DDPM dynamics (11) when initialized with $Y_{0}\sim X_{T}$;
  • $P^{\infty}$: the measure of DDPM dynamics (11) when initialized with $Y_{0}\sim\pi_{d}\overset{\mathrm{d}}{=}X_{\infty}$, where $X_{\infty}$ is the limit of the forward process (4) as $t\to\infty$.
@@
In this step, we list several lemmas that help connect the KL error of interest with the mean squared difference between the true and estimated posterior means. These results, which have been used in past work like Chen et al., 2022b ; Chen et al., 2023a ; Benton et al., (2024); Azangulov et al., (2024); Wu et al., (2024); Potaptchik et al., (2024), are now standard for analyzing DDPM-type samplers.
To begin with, we upper bound the KL error of interest, $\mathsf{KL}(q_{\delta}\,||\,p_{\mathsf{output}})$, by means of the KL divergence between two path measures. This is accomplished through the following standard lemma; its proof is fairly standard and we refer the interested reader to Benton et al., (2024, Section 3.3).
@@
If $\mathsf{KL}(Q\,||\,P^{T})<\infty$, then we have
$\mathsf{KL}(q_{\delta}\,||\,p_{\mathsf{output}})=\mathsf{KL}(q_{\delta}\,||\,p_{t_{N}})\leq\mathsf{KL}(Q\,||\,P^{\infty})=\mathsf{KL}(Q\,||\,P^{T})+\mathsf{KL}(q_{T}\,||\,\pi_{d}).$
In view of Lemma 3, to establish Theorem 2, it boils down to controlling $\mathsf{KL}(Q\,||\,P^{T})$ and $\mathsf{KL}(q_{T}\,||\,\pi_{d})$ separately. Let us start with the last term $\mathsf{KL}(q_{T}\,||\,\pi_{d})$, which is equivalent to analyzing the convergence rate of the forward process. The following lemma offers a simple upper bound on this term. Note that the exponential convergence of the OU process is well-known; we refer the reader to Chen et al., 2023a (, Lemma 9) or Benton et al., (2024, Appendix G) for the proof this lemma.
@@
As a result of Lemmas 3 and 4, it remains to control the term $\mathsf{KL}(Q\,||\,P^{T})$. The following lemma links $\mathsf{KL}(Q\,||\,P^{T})$ with the mean squared difference between the drift terms of SDEs (11b) and (13), a standard result that can be readily accomplished via the Girsanov theorem (Le Gall,, 2016).
@@
$\mathsf{KL}(Q\,||\,P^{T})\leq\sum_{n=0}^{N-1}\int_{t_{n}}^{t_{n+1}}\frac{1-\sigma^{2}_{T-t}}{\sigma_{T-t}^{4}}\mathbb{E}_{Q}\left[\big{\|}\mu_{T-t}(Y_{t})-\widehat{\mu}_{T-t_{n}}(Y_{t_{n}})\big{\|}_{2}^{2}\right]\mathrm{d}t<\infty.$ (20)
Proof.
The proof follows from an application of Girsanov’s Theorem to SDEs (11b) and (13), with the aid of a coupling argument. Given that this standard trick and its derivation has appeared in many recent papers (e.g., Chen et al., 2022b (, Section 5.2) or Benton et al., (2024, Appendix F)), we omit it for brevity. ∎
@@
In the above decomposition (21), $T_{1}$ reflects the influence of the score estimation error, whereas $T_{2}$ captures the effect of the time discretization error. In the sequel, we shall upper bound $T_{1}$ and $T_{2}$ separately, and combine them to reach our final bound on $\mathsf{KL}(Q\,||\,P^{T})$.
@@
To cope with the term $T_{1}$, we find it helpful to first establish the following result on the coefficients $\{\eta_{t}\}$, whose proof is provided in Appendix B.1. The arguments of this step are also fairly standard, akin to the ones used in, e.g., Azangulov et al., (2024); Potaptchik et al., (2024).
@@
Our approach shares similar spirits as Benton et al., (2024, Lemma 4). Similar characterizations were obtained in Section 4 of the concurrent work Potaptchik et al., (2024) by analyzing the martingale structure underlying the quantity of interest.
@@
Similar strategies were introduced in the previous work Benton et al., (2024, Appendix E), without accounting for the low-dimensional structure. It is worth noting that Lemma 9 in the concurrent work Potaptchik et al., (2024) shares similarity with the above result too. Note, however, that our lemma is established under much weaker assumptions on the low-dimensional structure as well as a much large range of the covariance of $X_{0}$, as discussed towards the end of Section 4.2.
@@
In the following analysis, we decompose the time horizon $[0,T-\delta]$ into $[0,T-1]$ and $[T-1,T-\delta]$, and control the right-hand side of equation (26) on these two time intervals separately. Without loss of generality, suppose that the $M$-th discretization time point obeys $t_{M}=T-1$.
@@
$\overset{\mathrm{(c)}}{\lesssim}\frac{(1-\sigma^{2}_{T-\kappa})\sigma^{2}_{T}}{1-\sigma^{2}_{T}}\kappa k\log k+\sum_{n=1}^{M-1}(1-\sigma_{T-{t_{n+1}}}^{2})\big{(}1-e^{-2(t_{n+1}-t_{n})}\big{)}\cdot\frac{\sigma^{2}_{T-t_{n}}}{1-\sigma^{2}_{T-t_{n}}}\kappa k\log k$
@@
$T_{2}$ $=\sum_{n=0}^{M-1}\int_{t_{n}}^{t_{n+1}}D_{t_{n},t}\mathrm{d}t+\sum_{n=M}^{N-1}\int_{t_{n}}^{t_{n+1}}D_{t_{n},t}\mathrm{d}t$
$\lesssim\kappa\min\left\{\mathbb{E}[\|X_{0}\|_{2}^{2}],Tk\log k\right\}+\kappa^{2}Nk\log k.$ (31)
@@
Combining inequalities (22) and (31) with Lemma 5, we obtain
@@ 6 Discussion
This paper has made progress towards demystifying the unreasonable effectiveness of then DDPM when the target distribution has a low-dimensional structure. We have sharpened the past result in Li and Yan, 2024a by proving that the iteration complexity scales linearly with the intrinsic dimension $k$ (note that linear dependency in $k$ was also established in the independent work Potaptchik et al., (2024), posted two weeks prior to our work). In some sense, our results reveal that the DDPM sampler is optimally adaptive to unknown low dimensionlity, provided that a suitable discretization schedule is adopted. Our theory leverages the intimate connection between the DDPM sampler and an adaptively discretized SDE, where the non-linear component of the drift time of this SDE is inherently low-dimensional.
We conclude this paper by pointing out several directions worthy of future investigation. To begin with, it remains unclear how the DDPM sampler adapts to unknown low dimensionality when other metrics (e.g., the total-variation distance or the Wasserstein distance) are used to measure distributional discrepancy. Another interesting direction is to examine the adaptivity of other mainstream samplers (e.g., the probability flow ODE (Song et al.,, 2021; Chen et al., 2023d, ; Li et al., 2024c, ; Gao and Zhu,, 2024; Huang et al.,, 2024)) in the face of unknown low-dimensional structure. Moreover, this paper treats the training procedure as a black box and directly assumes a mean squared score estimation error bound. For distributions with low intrinsic dimensions, a desirable score learning algorithm woulclud adapt to the low-dimensional structure too, which has been the focus of a recent line of work (Kadkhodaie et al.,, 2023; Tang and Yang,, 2024; Chen et al., 2024b, ; Azangulov et al.,, 2024; Chen et al., 2023b, ; Wang et al., 2024a, ; Wang et al., 2024b, ). It would thus be interesting to develop an end-to-end theory to accommodate both the pre-training phase and the sampling process. Finally, our analysis, which highlights the crucial role of proper parameterization and discretization, might shed light on how to design the coefficients for other algorithms (e.g., higher-order SDE and ODE solvers) in order to accommodate inherently structured data distributions.
@@ Acknowledgements
Y. Wei is supported in part by the NSF grant CCF-2418156 and the CAREER award DMS-2143215. Y. Chen is supported in part by the Alfred P. Sloan Research Fellowship, the AFOSR grant FA9550-22-1-0198, the ONR grant N00014-22-1-2354, and the NSF grant CCF-2221009. We would like to thank Jiadong Liang and Yuchen Wu for their inspiring discussions. We would also like to thank the authors of Potaptchik et al., (2024); Azangulov et al., (2024) for discussing the connections between our work and theirs, particularly for pointing out Azangulov et al., (2024, Proposition 28) that established the SDE equivalence of the DDPM.
@@ Appendix A Proof of Theorem 1
We proceed by directly solving SDE (11b). Let $(Y_{t})_{t\in[0,T-\delta]}$ be the solution to SDE (11b). Multiply $Y_{t}$ by some differentiable function $f(\cdot)$ (to be specified shortly) and apply Itô’s formula (Oksendal and Oksendal,, 2003) to yield
@@
With this choice of $f(\cdot)$ in place, one can eliminate the linear term in (33) and obtain the simplified form
@@
where we use the basic fact that $f(t)=\sqrt{1-\sigma_{T-t}^{2}}\sigma_{T-t}^{-2}$. Integrating both sides from $t_{n}$ to $t_{n+1}$ results in
@@
for some vector $\widetilde{Z}_{n}\sim\mathcal{N}(0,I_{d})$, where $\{\widetilde{Z}_{n}\}_{n=0,\dots,N-1}$ are statistically independent. This allows us to write
@@
By defining
$\gamma_{n}=e^{-(T-t_{n})},$
we can explicitly calculate the integral as follows
@@
=$\overset{\mathrm{(a)}}{=}\frac{1-\gamma_{n+1}^{2}}{\gamma_{n+1}}\cdot\frac{\gamma_{n}}{1-\gamma_{n}^{2}}Y_{t_{n}}+\frac{\gamma_{n+1}^{2}-\gamma_{n}^{2}}{(1-\gamma_{n}^{2})\gamma_{n}\gamma_{n+1}}\left(Y_{t_{n}}+(1-\gamma_{n}^{2})\widehat{s}_{T-t_{n}}(Y_{t_{n}})\right)+\sqrt{\frac{(\gamma_{n+1}^{2}-\gamma_{n}^{2})(1-\gamma_{n+1}^{2})}{(1-\gamma_{n}^{2})\gamma_{n+1}^{2}}}\widetilde{Z}_{n}$
@@
which coincides with the DDPM update rule (2) introduced in Ho et al., (2020).

### v3

@@ 5 Main analysis (Proof of Theorem 2)
We now present the proof of main convergence theory in Theorem 2. To control the KL divergence between the DDPM-type dynamics and our target distribution, a common approach is to connect the KL error with the $\ell_{2}$ discretization error by means of Girsanov’s theorem (e.g., Chen et al. (2022b, 2023b)). We shall follow this approach in this proof, while in the meantime taking into account the intrinsic low-dimensional structure of the target distribution. In particular, we shall invoke the covering argument to treat the low-dimensional structure approximately as the union of linear subspaces. Compared with previous work on low-dimensional adaptation (Azangulov et al., 2024; Li and Yan, 2024a), instead of focusing on the influence of the intrinsic dimension evolution of the score function, we improve the $k$-dependence by analyzing the posterior mean function, i.e., $\mu_{t}(\cdot)$. As it turns out, the evolution of the posterior mean function can be well calibrated along the forward process with high probability under this framework, which plays a key role in enabling our non-asymptotic convergence analysis.
Our main proof is composed of three steps, with the first two steps being fairly standard and the third step capturing the main challenges in the analysis of the posterior covariance matrix $\Sigma_{t}(x)$ (which is directly linked to the posterior mean function and discretization error). In particular, Lemma 10 reveals that the growth of the expected trace norm of $\Sigma_{t}(x)$ depends only on the intrinsic dimension $k$ as opposed to the ambient dimension $d$, thereby leading to a refined control of the discretization error. The analysis is supported by a sequence of auxiliary lemmas whose proofs are deferred to Appendices B and C. Throughout the proof, we adopt the following convenient notation:
@@
  • $P^{(T)}$: the measure of the DDPM dynamics (11) when initialized with $Y_{0}\sim X_{T}$;
  • $P^{(\infty)}$: the measure of DDPM dynamics (11) when initialized with $Y_{0}\sim\pi_{d}\overset{\mathrm{d}}{=}X_{\infty}$, where $X_{\infty}$ is the limit of the forward process (4) as $t\to\infty$.
@@
In this step, we list several lemmas that help connect the KL error of interest with the mean squared difference between the true and estimated posterior means. These results, which have been used in past work like Chen et al. (2022b, 2023a); Benton et al. (2024); Azangulov et al. (2024); Wu et al. (2024); Potaptchik et al. (2024), are now standard for analyzing DDPM-type samplers.
To begin with, we upper bound the KL error of interest, $\mathsf{KL}(q_{\delta}\,||\,p_{\mathsf{output}})$, by means of the KL divergence between two path measures $Q$ and $P^{(\infty)}$. This is accomplished through the following standard lemma; its proof is fairly standard and we refer the interested reader to Benton et al. (2024, Section 3.3).
@@
If $\mathsf{KL}(Q\,||\,P^{(T)})<\infty$, then we have
$\mathsf{KL}(q_{\delta}\,||\,p_{\mathsf{out}})=\mathsf{KL}(q_{\delta}\,||\,p_{t_{N}})\leq\mathsf{KL}(Q\,||\,P^{(\infty)})=\mathsf{KL}(Q\,||\,P^{(T)})+\mathsf{KL}(q_{T}\,||\,\pi_{d}).$
In view of Lemma 5, to establish Theorem 2, it boils down to controlling $\mathsf{KL}(Q\,||\,P^{(T)})$ and $\mathsf{KL}(q_{T}\,||\,\pi_{d})$ separately. Let us start with the last term $\mathsf{KL}(q_{T}\,||\,\pi_{d})$, which is equivalent to analyzing the convergence rate of the forward process. The following lemma offers a simple upper bound on this term. Note that the exponential convergence of the OU process is well-known; we refer the reader to Chen et al. (2023a, Lemma 9) or Benton et al. (2024, Appendix G) for the proof this lemma.
@@
As a result of Lemmas 5 and 6, it remains to control the term $\mathsf{KL}(Q\,||\,P^{(T)})$. The following lemma links $\mathsf{KL}(Q\,||\,P^{(T)})$ with the mean squared difference between the drift terms of SDEs (11b) and (12), a standard result that can be readily accomplished via the Girsanov theorem (Le Gall, 2016).
@@
$\mathsf{KL}(Q\,||\,P^{(T)})\leq\sum_{n=0}^{N-1}\int_{t_{n}}^{t_{n+1}}\frac{1-\sigma^{2}_{T-t}}{\sigma_{T-t}^{4}}\mathbb{E}_{Q}\left[\big\|\mu_{T-t}(Y_{t})-\widehat{\mu}_{T-t_{n}}(Y_{t_{n}})\big\|_{2}^{2}\right]\mathrm{d}t<\infty.$ (19)
The proof follows from an application of Girsanov’s Theorem to SDEs (11b) and (12), with the aid of a coupling argument. Given that this standard trick and its derivation has appeared in many recent papers (e.g., Chen et al. (2022b, Section 5.2) or Benton et al. (2024, Appendix F)), we omit it for brevity.
@@
In the above decomposition (20), $T_{1}$ reflects the influence of the score estimation error, whereas $T_{2}$ captures the effect of the time discretization error. In the sequel, we shall upper bound $T_{1}$ and $T_{2}$ separately, and combine them to reach our final bound on $\mathsf{KL}(Q\,||\,P^{(T)})$.
@@
To cope with the term $T_{1}$, we find it helpful to first establish the following result on the coefficients $\{\eta_{t}\}$, whose proof is provided in Appendix B.1. The arguments of this step are also fairly standard, akin to the ones used in, e.g., Azangulov et al. (2024); Potaptchik et al. (2024).
@@
Our approach shares similar spirits as Benton et al. (2024, Lemma 4). Similar characterizations were obtained in Section 4 of the concurrent work Potaptchik et al. (2024) by analyzing the martingale structure underlying the quantity of interest.
@@
Similar strategies were introduced in the previous work Benton et al. (2024, Appendix E), without accounting for the low-dimensional structure. It is worth noting that Lemma 9 in the concurrent work Potaptchik et al. (2024) shares similarity with the above result too. Note, however, that our lemma is established under much weaker assumptions on the low-dimensional structure as well as a much large range of the covariance of $X_{0}$, as discussed towards the end of Section 4.2.
@@
In the following analysis, we decompose the time horizon $0,T-\delta]$ into [0,T1][0,T-1] and $T-1,T-\delta]$, and control the right-hand side of equation (25) on these two time intervals separately. Without loss of generality, suppose that the $M$-th discretization time point obeys $t_{M}=T-1$.
@@
$\overset{\mathrm{(c)}}{\lesssim}\frac{(1-\sigma^{2}_{T-\kappa})\sigma^{2}_{T}}{1-\sigma^{2}_{T}}\kappa k\log k$
$\qquad\qquad\qquad+\sum_{n=1}^{M-1}(1-\sigma_{T-{t_{n+1}}}^{2})\big(1-e^{-2(t_{n+1}-t_{n})}\big)\cdot\frac{\sigma^{2}_{T-t_{n}}}{1-\sigma^{2}_{T-t_{n}}}\kappa k\log k$
@@
$T_{2}$ $=\sum_{n=0}^{M-1}\int_{t_{n}}^{t_{n+1}}D_{t_{n},t}\mathrm{d}t+\sum_{n=M}^{N-1}\int_{t_{n}}^{t_{n+1}}D_{t_{n},t}\mathrm{d}t\lesssim\kappa\min\left\{\mathbb{E}[\|X_{0}\|_{2}^{2}],Tk\log k\right\}+\kappa^{2}Nk\log k.$ (30)
@@
Combining inequalities (21) and (30) with Lemma 7 gives
@@ 6 Discussion
This paper has made progress towards demystifying the unreasonable effectiveness of then DDPM when the target distribution has a low-dimensional structure. We have sharpened the past result in Li and Yan (2024a) by proving that the iteration complexity scales linearly with the intrinsic dimension $k$ (note that linear dependency in $k$ was also established in the concurrent work Potaptchik et al. (2024)). In some sense, our results reveal that the DDPM sampler is optimally adaptive to unknown low dimensionlity, provided that a suitable discretization schedule is adopted. Our theory leverages the intimate connection between the DDPM sampler and an adaptively discretized SDE, where the non-linear component of the drift time of this SDE is inherently low-dimensional.
We conclude this paper by pointing out several directions worthy of future investigation. To begin with, it remains unclear how the DDPM sampler adapts to unknown low dimensionality when other metrics (e.g., the total-variation distance or the Wasserstein distance) are used to measure distributional discrepancy. For example, the FID score — commonly adopted in practice — is not fully captured by the KL divergence, but instead should be interpreted as a mixture of other metrics. However, existing convergence analysis under other metrics typically requires much more stringent assumptions like smoothness on the (noisy) score function and/or log-concavity on the target distribution, which remains much less developed. Another interesting direction is to examine the adaptivity of other mainstream samplers (e.g., the probability flow ODE (Song et al., 2021; Chen et al., 2023d; Li et al., 2024c; Gao and Zhu, 2024; Huang et al., 2024; Liang et al., 2025; Li et al., 2025c)) in the face of unknown low-dimensional structure. Although the ODE-based samplers often achieve better iteration complexity than the SDE counterpart in practice, the convergence rate of ODE-based sampler in KL divergence indeed cannot surpass the linear rate. It would be great to see whether ODE samplers can achieve more favorable convergence guarantees under certain practically relevant metrics like the Wasserstein distance. Moreover, this paper treats the training procedure as a black box and directly assumes a mean squared score estimation error bound. For distributions with low intrinsic dimensions, a desirable score learning algorithm would adapt to the low-dimensional structure too, which has been the focus of a recent line of work (Kadkhodaie et al., 2023; Tang and Yang, 2024; Chen et al., 2024b; Azangulov et al., 2024; Chen et al., 2023b; Wang et al., 2024a, b). It would thus be interesting to develop an end-to-end theory to accommodate both the pre-training phase and the sampling process. Finally, our analysis, which highlights the crucial role of proper parameterization and discretization, might shed light on how to design the coefficients for other algorithms (e.g., higher-order SDE and ODE solvers) in order to accommodate inherently structured data distributions.
@@ Acknowledgements
Y. Wei is supported in part by the NSF grant CCF-2418156 and the CAREER award DMS-2143215. Y. Chen is supported in part by the Alfred P. Sloan Research Fellowship, the ONR grants N00014-22-1-2354 and N00014-25-1-2344, the NSF grants 2221009 and 2218773, the Wharton AI & Analytics Initiative’s AI Research Fund, and the Amazon Research Award. We would like to thank Jiadong Liang and Yuchen Wu for their inspiring discussions. We would also like to thank the authors of Potaptchik et al. (2024); Azangulov et al. (2024) for discussing the connections between our work and theirs, particularly for pointing out Azangulov et al. (2024, Proposition 28) that established the SDE equivalence of the DDPM.
@@ Appendix A Proof of Theorem 1
We proceed by directly solving SDE (11b). Let $(Y_{t})_{t\in[0,T-\delta]}$ be the solution to SDE (11b). Multiply $Y_{t}$ by some differentiable function $f(\cdot)$ (to be specified shortly) and apply Itô’s formula (Oksendal and Oksendal, 2003) to yield
@@
With this choice of $f(\cdot)$ in place, one can eliminate the linear term in (32) and obtain
@@
where we use the basic fact that $f(t)=\sqrt{1-\sigma_{T-t}^{2}}\sigma_{T-t}^{-2}$. Integrating both sides from $t_{n}$ to $t_{n+1}$ gives
@@
for some vector $\widetilde{Z}_{n}\sim\mathcal{N}(0,I_{d})$, where $\{\widetilde{Z}_{n}\}_{n=0,\dots,N-1}$ are independent. This allows us to write
@@
By defining $\gamma_{n}=e^{-(T-t_{n})}$. we can explicitly calculate the integral as follows
@@
$\overset{\mathrm{(a)}}{=}\frac{1-\gamma_{n+1}^{2}}{\gamma_{n+1}}\cdot\frac{\gamma_{n}}{1-\gamma_{n}^{2}}Y_{t_{n}}+\frac{\gamma_{n+1}^{2}-\gamma_{n}^{2}}{(1-\gamma_{n}^{2})\gamma_{n}\gamma_{n+1}}\left(Y_{t_{n}}+(1-\gamma_{n}^{2})\widehat{s}_{T-t_{n}}(Y_{t_{n}})\right)$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\sqrt{\frac{(\gamma_{n+1}^{2}-\gamma_{n}^{2})(1-\gamma_{n+1}^{2})}{(1-\gamma_{n}^{2})\gamma_{n+1}^{2}}}\widetilde{Z}_{n}$
@@
which coincides with the DDPM update rule (2) introduced in Ho et al. (2020).